a telephone poll of 758 Arizona adults who celebrate Christmas. Question asked was do you open presents on Christmas Eve or Christmas Day. Of those surveyed, 394 said they wait until Christmas Day.
a. Determine the sample proportion.
b. At the 5% significance level, do the data provide sufficient evidence to conclude that the majority of Arizona families who celebrate Christmas wait until Christmas Day to open their presents?
Any help you can give me will be very appreciated.

Answer by stanbon(75887)

a telephone poll of 758 Arizona adults who celebrate Christmas. Question asked was "Do you open presents on Christmas Eve or Christmas Day?". Of those surveyed, 394 said they wait until Christmas Day.
a. Determine the sample proportion.:::394/758 = 0.5198
-----------------------------
b. At the 5% significance level, do the data provide sufficient evidence to conclude that the majority of Arizona families who celebrate Christmas wait until Christmas Day to open their presents?
----
Ho: p <= 1/2
Ha: p > 1/2
-----
p-hat = 0.5198
test statistic: z(0.5198) = (0.5198-0.5)/sqrt[(1/2)(1/2)/758] = 2.1805
----
p-value: P(z> 2.1805) = 0.0146
----------------
Conclusion: Since the p-value is less than 5%, reject Ho.
The test results do not support Ho.
There is some evidence that most wait till Christmas Day.
==========================================
Cheers,
Stan H.
========